1) Congratulations! The FDA has just approved your company's new drug, and now ABC Pharmaceuticals wants to buy its North American marketing rights. Before you decide whether to sell, you must assess the four payment plan options offered by ABC Pharmaceuticals:

  • Option A: Receive annual payments of $25 million forever.

  • Option B: Receive annual payments forever starting at $15 million and increasing each year with the inflation rate of 3% per year.

  • Option C: Receive 25 annual payments of $27.5 million.

  • Option D: Receive 25 annual payments starting at $22.5 million and increasing each year with the inflation rate of 3% per year.

What is the present value of each option? For each payment plan option, assume the first payment would begin one year from today, and that the nominal discount rate is 10% per year. (Note: Your answer should be expressed in units of millions of dollars.)

NPV Option A = $____million

NPV Option B = $____million

NPV Option C = $____million

NPV Option D = $____million

For Option D, note that the present value of a growing annuity that pays ? dollars starting one year from now with subsequent annual payments that grow at a rate ? is given by,

??(?)=??−?×[1−(1+?1+?)?]

where ? is the discount rate and ? is the time until the last payment. Optional: Can you derive this formula? Think about subtracting a growing perpetuity that starts paying out in year ?+1 from a growing perpetuity that starts paying out 1 year from today.

2) Consider the medication Sovaldi. Marketed by Gilead Sciences for the treatment of hepatitis C, Sovaldi has cure rates of 30% to 97%, depending on the type of hepatitis C virus. In 2014, soon after the drug was approved, the list price for a 12-week course of treatment in the US was $84,000. For many representatives in Congress, this was an outrageous price, causing the U.S. Senate Finance Committee to conduct an eighteen-month investigation into Gilead’s pricing strategy.

Now consider buying a one-time cure for a debilitating disease like hepatitis C in the same way a prospective car owner buys a car. Just as a Tesla provides the car owner with a perpetual stream of services to drive that car, a cure can be thought of as a perpetual stream of health benefits for the patient to live in a healthy body. From this perspective, we can match the stream of payments with the stream of health benefits, turning the cost into a much more manageable monthly mortgage. Given a set loan term of 10 years and a fixed monthly interest rate of 0.4167%, calculate the monthly payment on a cure that costs $84,000. (Note: Your answer should be expressed to 2 decimal places.)

Monthly payment = $_____
Answer #1

                 

  • Option A: Use the perpetuity formula.

    PV=Cr=250.1=$250 million

  • Option B: Use the growing perpetuity formula.

    PV=Crg=150.100.03=$214.29 million

  • Option C: Use the annuity formula.

    PV=Cr(11(1+r)T)=27.50.10(11(1+0.10)25)=$249.62 million

  • Option D: Use the growing annuity formula.

    PV=Crg(1(1+g1+r)T)=22.50.100.03(1(1+0.031+0.10)25)=$259.31 million

Option D has the largest present value.

         

 

        

 


answered by: palak garg